OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (anonymous):

@luckythebest how do we do this ?

OpenStudy (anonymous):

What is formula of slope?

OpenStudy (anonymous):

Slope = y2-y1/x2-x1.

OpenStudy (anonymous):

ok. and slope given is -3/2 so -3/2 = y2-y1/x2-x1. try and solve. :)

OpenStudy (anonymous):

what does y and x =

OpenStudy (anonymous):

x1 = x, y1 = 6, x2 = 10, y2 = -3 and -3/2 = y2-y1/x2-x1 Solve for x.

OpenStudy (anonymous):

i dont understand :( but ill try

OpenStudy (anonymous):

-3/2 = -3-6/10-x

OpenStudy (anonymous):

-3-6=-9

OpenStudy (anonymous):

yep. -3/2 = -9/10-x -3(10-x) = 2.-9 at least now? :P

OpenStudy (anonymous):

so x = 2.9 ? how did you get 2.9 ?

OpenStudy (anonymous):

^ is the equation. -3 x (10-x) = 2 x -9 i.e. -30 + 3x = -18

OpenStudy (anonymous):

that what i got was 30 that why i asked you how you got 2.9

OpenStudy (anonymous):

ok ok i get it. :P 2.-9 i meant 2 x -9 ( . means multiplication)

OpenStudy (anonymous):

so now we have -9/-18 ?

OpenStudy (anonymous):

Ok wait here. -3/2 = -9/10-x -3(10-x) = 2(-9) => -30 + 3x = -18 => 3x = 12 i.e. x = 4.

OpenStudy (anonymous):

ooooooooooooooooooo.